The sum of two numbers is 16. Three times the larger number decreased by the smaller number is 72. Find the two numbers.
Answer by macston(5194)

x=the larger number; y=the smaller number
x+y=16
x=16-y
3x-y=72 Substitute for x
3(16-y)-y=72
48-3y-y=72 Subtract 48 from each side.
-4y=24 Divide each side by -4
y=-6 ANSWER 1:The smaller number is -6.
x=16-y=16-(-6)=22 ANSWER 2: The larger number is 22.
CHECK:
3x-y=72
3(22)-(-6)=72
66-(-6)=72
72=72